OpenStudy (anonymous):

Solve 2x2 − 8x = −7. negative 2 plus or minus square root of 2 negative 2 plus or minus 2 square root of 2 quantity of 2 plus or minus square root of 2 all over 2 2 plus or minus square root of 2 end root over 2

OpenStudy (anonymous):

correct answer will get medal and fan

OpenStudy (anonymous):

@aaronq @Nnesha

Nnesha (nnesha):

set it equal to zero

Nnesha (nnesha):

then use quadratic\[\huge\rm x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] formula a b and c are \[\large\rm Ax^2+Bx+C=0\] a=leading coefficient b=middle term c=constant term

OpenStudy (anonymous):

a=3,b=−7,c=3

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

bo

Nnesha (nnesha):

\[2x^2-8x=-7\] first set it equal to zero write it in Ax^2+Bx+C=0 form add 7 both sides

OpenStudy (anonymous):

so b is correct @Nnesha

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha bo \(\color{blue}{\text{End of Quote}}\) correction no

Nnesha (nnesha):

add 7 both sides let me know what you get

OpenStudy (anonymous):

4 ± 2√ 2

OpenStudy (anonymous):

but thats not an answer choice @Nnesha

Nnesha (nnesha):

that should be in fraction form

Nnesha (nnesha):

a= ?? there is 2a at the denominator replace a by its value and multiply that with 2

Nnesha (nnesha):

check your equation is there a typo somwhere ?